OpenStudy (anonymous):

Hello I just needed someone's help on my PRE-ALG Questions 2(x-3) = 2x 1.one solution 2.no solution 3. infinitely many solutions 2. 3(y - 3) = 2y - 9 + y 1. one solution 2. no solutions 3. infinitely many solutions 3. 10x - 2 - 6x = 3x - 2 + x 1. one solution 2. no solutions 3. infinitely many solutions 4. 4(x + 3) + 2x = x - 8 1. one solution 2. no solution 3. infinitely many solutions

OpenStudy (anonymous):

Hi what exactly are you stuck with?

OpenStudy (anonymous):

I have problem with understanding it

OpenStudy (anonymous):

Ok, so do you know what "having a solution" is?

OpenStudy (anonymous):

yeah

OpenStudy (jidnyesha):

the first 3 have no solutions n the last has one solution coz in the first 3 equations when u solve it the x gets cancelled

OpenStudy (jidnyesha):

when u simplify the first equation both sides have 2x so it gets cancelled same wid the next 2 equations but only in the last one u get one solution

OpenStudy (anonymous):

the first one has no solution, while the second and the third one have infinite amount of solutions. The last one has one solution.

OpenStudy (anonymous):

o ok i see

OpenStudy (anonymous):

The reason for the first one having no solution is because when you divide both sides by two: \[\frac{2(x-3) }{ 2 }=\frac{ 2x }{2 }\] you get x-3 = x which we know is false, because x-3 does not equal x. Therefore there is no solution for this.

OpenStudy (jidnyesha):

wait y the second and third have infinite solutions?

OpenStudy (anonymous):

o i see now i was doing it wrong the whole time lol thank you on my other tests

OpenStudy (anonymous):

However, for the second and third one, through simplification of the equations, we have: 2) 3y-9 = 3y -9 3) 4x-2 = 4x-2 Notice how the functions on both sides of the equal sign are the same. This means that both functions will overlap each other entirely when graphed. They will always equal to each other.

OpenStudy (jidnyesha):

i took the first one as \[2x-6=2x\] \[-6=0\] which isnt possible so 0 solutions

OpenStudy (jidnyesha):

i still didnt get the reason for second and third equations having infinite solutions

OpenStudy (anonymous):

Ok. Looking at this graphically, "having a solution" means that both functions will cross each other (aka the values of both functions will equal at some point). Having one solution means the value of both functions will equal at one point. Having no solution means the values of both functions will never equal. Having an infinite amount of solutions means that the values of both functions will always equal to each other. Given that for #2 and #3, we have an identical function on each side of the equal sign, they will always equal to each other. So they will always overlap each other graphically.

OpenStudy (anonymous):

thanks for both of your help.

OpenStudy (jidnyesha):

oh thnx for explaining

OpenStudy (anonymous):

no probs